# Some Optimizers are More Equal: Understanding the Role of Optimizers in Group Fairness

### v1

@@ Abstract
We study whether and how the choice of optimization algorithm can impact group fairness in deep neural networks. Through stochastic differential equation analysis of optimization dynamics in an analytically tractable setup, we demonstrate that the choice of optimization algorithm indeed influences fairness outcomes, particularly under severe imbalance. Furthermore, we show that when comparing two categories of optimizers, adaptive methods and stochastic methods, RMSProp (from the adaptive category) has a higher likelihood of converging to fairer minima than SGD (from the stochastic category). Building on this insight, we derive two new theoretical guarantees showing that, under appropriate conditions, RMSProp exhibits lower parameter bias and improved fairness in a single optimization step compared to SGD. We then validate these findings through extensive experiments on three publicly available datasets, namely CelebA, FairFace, and MS-COCO, across different tasks as facial expression recognition, gender classification, and multi-label classification, using various backbones. Considering multiple fairness definitions including equalized odds, equal opportunity, and demographic parity, adaptive optimizers like RMSProp and Adam consistently outperform SGD in terms of group fairness, while maintaining comparable predictive accuracy. Our results highlight the role of adaptive updates as a crucial yet overlooked mechanism for promoting fair outcomes. We release the source code at: https://anonymous.4open.science/??????.
@@ 1 Introduction
Machine learning is increasingly applied to a wide variety of domains, many of which have significant social implications [pmlr-v81-buolamwini18a, 10.1145/3494672]. These include, but are not limited to, decision-making systems, risk assessment models, recommendation engines, and personalized health interventions [HEGDE2020104492, rahimi2022quantum]. These systems influence people’s lives and can lead to unintended biases or disparities [10.1145/3457607], for instance by treating different groups of people differently. To address this, researchers have developed various techniques to ensure group fairness in machine learning algorithms [10.1145/3577190.3614141, hosseini2025faces].
Prior work has demonstrated that a variety of factors such as training data imbalance [10.1007/978-3-030-87199-4_39] and model hyper-parameters [NEURIPS2023_eb3c42dd] can impact the performance of deep neural networks in terms of group fairness and bias. In this paper, we pose the following question for the first time:
@@
In this paper, we bridge this gap by first analyzing the impact of optimization in group fairness in a toy setup. This initial study confirms our hypothesis: that indeed the choice of optimizer impacts group fairness in downstream applications, particularly in the presence of severe bias in the training set. Next, we delve deep into the problem by proposing and proving two new theorems that demonstrate that among the two primary families of optimization algorithms, adaptive gradient methods such as Adam and RMSProp, and stochastic gradient methods such as SGD, the former is more likely to exhibit better group fairness. Next, we validate our theoretical findings by conducting extensive experiments on three publicly available datasets-MS-COCO [10.1007/978-3-319-10602-1_48], CelebA [liu2015faceattributes], and FairFace [karkkainenfairface] across three different tasks: multi-label classification, facial expression recognition, and gender classification. We carefully evaluate fairness using three widely adopted metrics: Equalized Odds, Equal Opportunity, and Demographic Parity. The results are strongly aligned with our theoretical analysis, consistently demonstrating that adaptive gradient methods achieve better fairness scores compared to stochastic gradient methods.
Our contributions in this paper are summarized as follows: (1) First, we provide a detailed mathematical analysis to study the impact of optimization algorithms on fairness in a toy setting. Here, we demonstrate that adaptive gradient algorithms like RMSProp have a higher probability of converging to fairer minima. (2) Next, we establish and prove two theorems showing that under appropriate conditions: (a) RMSProp provides fairer updates than SGD, and (b) the worst-case bias introduced by RMSProp in a single optimization step has an upper bound equal to that of SGD. These theoretical insights highlight key properties of optimization algorithms that contribute to their advantages in terms of fairness. (3) Through extensive experiments on multiple publicly available datasets and tasks, we empirically validate our theoretical findings. We systematically evaluate the fairness of models optimized using adaptive gradient methods and compare the outcomes to those obtained from stochastic gradient methods. Our results consistently show that adaptive gradient methods lead to fairer minima. (4) To enable rapid reproducibility and contribute to the area, we make our code publicly available at: https://anonymous.4open.science
@@ 2 Related work
Group fairness. Bias in machine learning models lead to discriminatory outcomes, and as these models are increasingly being deployed in real-life applications, such biases can imposes risk to the society [10.1145/3457607]. This risk is particularly important in high-stake applications, as hiring [10.1093/ojls/gqab006], healthcare [lett2023translating], etc. When models yield biased outcomes, they favor certain demographic groups, which reinforces the current social inequalities. As an example, facial expression recognition models show higher error rates for darker skin tones [10.1007/978-3-030-65414-6_35]. Such biases undermine trust in machine learning models and pose ethical as well as legal challenges. To analyze fairness in machine learning, researchers in this area often distinguish the notion of “group fairness” and “individual fairness” [10.1145/3457607]. Group fairness focuses on this notion that machine learning models treat different demographic groups, such as those defined by race, the same. On the other hand, individual fairness requires that similar individuals be treated similarly. We focus on group fairness in this paper, as it is more recognized in practical applications [10.1145/3630106.3659006].
@@
One of the main sources of bias in machine learning models is data imbalance [deng2023fifa]. When certain demographic subgroups are underrepresented, models prioritize patterns from the majority group, leading to biased decision. To tackle this, a variety of methods have been proposed that can be categorized into 3 main groups [pmlr-v162-chai22a]: pre-process, in-process, and post-process. pre-process methods often apply some type of transformation to data to make the demographic information less recognizable to the model. As an example, [10.1145/3375627.3375865] proposes to augment the train set by synthetically generated samples, thereby ensuring an equal distribution for demographic subgroups. In-process methods directly modify the lerarning algorithm of the model. As an example, [10.1145/3577190.3614141] proposes a new loss function based of maximum mean discrepancy statistical distance to enhance fairness. Post-process methods intervenes after training is finished. For instance,  [NEURIPS2021_d9fea4ca] introduces a post-processing approach using graph Laplacian regularization to enforce individual fairness constraints while preserving accuracy. Nevertheless, in practice, these fairness enhancing methods are rarely used in real-world pipelines [Olulana_Cachel_Murai_Rundensteiner_2024]. One major limitation is their computational overhead or interference with the training phase. As an example, pre-processing methods require extensive data augmentation or transformations. Considering these challenges, an alternative approach to enhance fairness is to use components that are inherently present in every training. One of the fundamental building blocks of modern deep learning is optimization [NEURIPS2024_646d2edf]. Understanding the role of different optimization algorithms in fairness is crucial for practitioners seeking to develop fairer models without additional fairness-specific modifications.
@@
Optimization algorithms. Optimization algorithms in deep learning generally fall into two main categories: SGD and adaptive gradient methods, such as Adam [kingma2014adam] and RMSProp [maunderstanding]. SGD updates model parameters using a constant learning rate [NEURIPS2024_ca98452d], while adaptive methods dynamically adjust per-parameter learning rates based on past gradients, often leading to faster convergence [ma2022qualitative].The impact of these 2 families of optimization algorithms on group fairness remains less understood. Nonetheless, several studies have explored their effects in other areas. For instance, [maunderstanding] empirically demonstrate that models trained with SGD are more robust to input perturbations than those trained with adaptive methods. Fairness and robustness are often competing objectives in deep learning [NEURIPS2022_a80ebbb4]. This hints that adaptive optimizers may yield better fairness compared to SGD. However, this relationship has been neither studied, nor understood.
@@ 3.1 Preliminaries
Optimization algorithms. Stochastic gradient methods like SGD and adaptive gradient methods such as RMSProp and Adam are widely used for optimizing deep neural networks [NEURIPS2020_08fb104b]. Given a loss function $\mathcal{L}(w)$, SGD updates the model parameters $w$ at each iteration $k$ as:
@@
where $\eta$ denotes the learning rate. Adaptive gradient methods improve the speed of convergence by applying coordinate-wise scaling to the gradient [pan2022toward]. Specifically, RMSProp normalizes the gradient with an exponentially decaying average of its squared values:
@@
Stochastic Differential Equations (SDE). Given the interval $[0,T]$ where $T>0$, we have  [song2021scorebased]:
@@
where $X_{t}\in R^{d}$ is a stochastic process (the solution of the SDE), $a:R^{d}\times[0,T]\rightarrow R^{d}$ is the drift term, $b:R^{d}\times[0,T]\rightarrow R^{d\times s}$ is the diffusion term, and $B(t)\in R^{s}$ is $s$-dimensional Brownian motion. Intuitively, SDEs extend ordinary differential equations by incorporating stochastic noise. Analyzing functions involving stochastic processes requires tools beyond classical calculus. A fundamental theorem in stochastic analysis is Itô’s lemma, extending the chain rule to the functions of stochastic processes [anresampling]. Please see the Appendix A for Itô’s lemma. Unlike ODEs, the solution of an SDE is a stochastic process $X_{t}$. Formally, the solution can be represented as:
[…]
@@
[…]
where $\pi(n)$ denotes a sequence of $n$ partitions in $[0,t]$, and $\lim$ shows convergence in probability. In many cases, explicitly computing $X_{t}$ is intractable. Instead, it is often more practical to analyze the probability distribution of $X_{t}$. The Fokker-Planck equation, also known as the forward Kolmogorov equation, describes the time evolution of the probability density function $p(x,t)$ of $X_{t}$ [lai2023fp] as:
@@
[…]
SDE approximations for optimization algorithms. SDEs provide a powerful framework for analyzing the dynamics of discrete optimization algorithms such as SGD and RMSProp in a continuous setup. While direct analysis of discrete updates is often challenging [NEURIPS2021_69f62956], SDE approximations enable the study of optimization trajectories. In the context of optimization, the sources of stochasticity include mini-batch gradient noise which are variations due to data heterogeneity (e.g., differences in mini-batches drawn from different subpopulations), among others. To formalize this, we adopt the Noisy Gradient Oracle with Scale Parameter (NGOS) [NEURIPS2022_32ac7101], which models stochastic gradients as $g(w)=\nabla\mathcal{L}(w)+\Theta z$, where $\Theta$ is a noise scale parameter, and $z$ is a random variable that follows a distribution with covariance matrix $\Sigma(w)$ and mean of zero. This captures the inherent randomness in stochastic optimization. A well-behaved NGOS is defined as $\nabla\mathcal{L}(w)$ being Lipschitz and $\sqrt{\Sigma(w)}$ being bounded. See Appendix C for a detailed description of a well-behaved NGOS. Note that these assumptions are common and standard in prior works that study optimization using SDEs [li2017stochastic].
Under well-behaved NGOS, SGD can be approximated by the following SDE [li2017stochastic]:
@@
where $W_{t}$ represents an approximation of $w_{k}$ at discrete time steps $k\,\eta$. More recently, SDE approximation for RMSProp has been proposed as a coupled system of equations [NEURIPS2022_32ac7101]:
@@ 3.2 SDE analysis of optimizers
where the weights correspond to the proportion of each subgroup in the population. In this setup, the optimal solution for subgroup 0 is $w_{0}^{*}=1$, while the optimal solution for subgroup 1 is $w_{1}^{*}=-1$. However, The population’s optimal solution is obtained by minimizing $\mathcal{L}_{\text{pop}}(w)$, leading to $w_{\text{pop}}^{*}=0$. The following lemma shows that under the demographic parity definition of group fairness [10.1145/3457607], $w_{\text{pop}}^{*}=0$ is the fairest minima.
@@
In practice, we do not have access to the full population above, rather we have access to the set $\Omega$ of $N$ individual samples, sampled uniformly from the population. Additionally, set $\Omega$ may not contain an equal number of samples from subgroups 0 and 1, thus introducing bias. We assume that the probability of subgroup 0 in $\Omega$ is $p_{0}$ while the probability of subgroup 1 is $p_{1}=1-p_{0}$. For a given set $\Omega$, the empirical minimization problem used in practice is defined as: $\mathcal{L}_{emp}(w)=\frac{1}{N}\sum_{r\in\Omega}\mathcal{L}_{q_{r}}(w),$ where $q_{r}\in\{0,1\}$ corresponds to samples from subgroups 0 and 1. Note that as $N\rightarrow\infty$, $\mathcal{L}_{emp}\rightarrow\mathcal{L}_{pop}$. In min-batch training with batch size of 1, at each iteration, the optimizer processes a single individual from subgroup 0 with a probability of $p_{0}$ and from subgroup 1 with a probability $p_{1}$. In other words, with both SGD and RMSProp, the model receives $\nabla\mathcal{L}_{0}(w)$ with a probability of $p_{0}$ and $\nabla\mathcal{L}_{1}(w)$ with a probability of $p_{1}$, updating $w$ according to Eqs. 1 and 2, respectively. If $p_{0}>p_{1}$, the optimization trajectory is biased towards the group-specific minimum at $w=1$, and $w=-1$ otherwise. This demonstrates that disproportionate sampling of subgroups leads to unfair minima, which is particularly relevant in practical settings. The following theorem establishes that when the sampling bias $|p_{0}-p_{1}|$ exceeds a certain threshold, RMSProp has a higher probability of converging to $w^{*}_{pop}$ compared to SGD.
@@
Here, $\epsilon$ is added for numerical stability, but it can be ignored as it is commonly done so in prior work [NEURIPS2022_32ac7101]. Next, let’s derive the SDE for SGD:
@@
where $\vartheta=\frac{1}{8\eta p_{0}p_{1}}$. As we use a batch size of 1, we can set $\Theta=1$ [NEURIPS2022_32ac7101]. Then, $1>=2\sqrt{p_{0}p_{1}}$ inequality holds. With this in mind, by substituting $w^{*}_{pop}=0$ in Eqs. 20 and 21, we can conclude $\frac{p_{rms}(w^{*}_{pop})}{p_{sgd}(w^{*}_{pop})}>1$ holds if and only if $|p_{0}-p_{1}|>\Delta(p_{1}p_{2},\eta)$, where $\Delta(p_{1}p_{2},\eta)=\sqrt{\frac{\frac{1}{2}\ln{\frac{\vartheta}{\kappa}}}{\vartheta-\kappa}}$. ∎
@@
Now, let’s perform a quick simulation to verify our findings. We use the loss functions defined in Eq. 13 and set $p_{0}=0.1$ and $p_{1}=0.9$ to induce a severe sampling bias. We fix the learning rate to $\eta=0.1$ and apply SGD and RMSProp for 100 epochs. Each experiment is repeated 1000 times, and at every epoch we record the fraction of runs converging to the neighborhood of the fair optimum $w^{*}_{pop}$, defined by $|w-w^{*}_{pop}|<0.2$. The result is shown in Fig. 1 (left). As it is evident from this figure, approximately 10% of the runs under RMSProp lie in the neighborhood of $w^{*}_{pop}$, whereas SGD rarely converges to this fair region. Now, let’s induce a milder bias by setting $p_{0}=0.3$ and $p_{1}=0.7$ and run the same experiment. The results are illustrated in Fig. 1 (right). As shown in this figure, in the milder bias, the two optimizers behave more similarly, yet RMSProp attains a slightly higher fraction of solutions near $w^{*}_{pop}$. These observations align with our theoretical findings in Theorem 1. To show that our findings are true for a wider range of $\eta$ and alternative definitions of the fair neighborhood, we refer the reader to Appendix H for additional experiments.

### v2

@@ Abstract
We study whether and how the choice of optimization algorithm can impact group fairness in deep neural networks. Through stochastic differential equation analysis of optimization dynamics in an analytically tractable setup, we demonstrate that the choice of optimization algorithm indeed influences fairness outcomes, particularly under severe imbalance. Furthermore, we show that when comparing two categories of optimizers, adaptive methods and stochastic methods, RMSProp (from the adaptive category) has a higher likelihood of converging to fairer minima than SGD (from the stochastic category). Building on this insight, we derive two new theoretical guarantees showing that, under appropriate conditions, RMSProp exhibits fairer parameter updates and improved fairness in a single optimization step compared to SGD. We then validate these findings through extensive experiments on three publicly available datasets, namely CelebA, FairFace, and MS-COCO, across different tasks as facial expression recognition, gender classification, and multi-label classification, using various backbones. Considering multiple fairness definitions including equalized odds, equal opportunity, and demographic parity, adaptive optimizers like RMSProp and Adam consistently outperform SGD in terms of group fairness, while maintaining comparable predictive accuracy. Our results highlight the role of adaptive updates as a crucial yet overlooked mechanism for promoting fair outcomes. We release the source code at: https://github.com/Mkolahdoozi/Some-Optimizers-Are-More-Equal.
@@ 1 Introduction
Machine learning is increasingly applied to a wide variety of domains, many of which have significant social implications [5, 43]. These include, but are not limited to, decision-making systems, risk assessment models, recommendation engines, and personalized health interventions [12, 46]. These systems influence people’s lives and can lead to unintended biases or disparities [38], for instance by treating different groups of people differently. To address this, researchers have developed various techniques to ensure group fairness in machine learning algorithms [20, 13].
Prior work has demonstrated that a variety of factors such as training data imbalance [45] and model hyper-parameters [10] can impact the performance of deep neural networks in terms of group fairness and bias. In this paper, we pose the following question for the first time:
@@
In this paper, we bridge this gap by first analyzing the impact of optimization in group fairness in an analytically tractable setup. This initial study confirms our hypothesis: that indeed the choice of optimizer impacts group fairness in downstream applications, particularly in the presence of severe bias in the training set. Next, we delve deep into the problem by proposing and proving two new theorems that demonstrate that among the two primary families of optimization algorithms, adaptive gradient methods such as Adam and RMSProp, and stochastic gradient methods such as SGD, the former is more likely to exhibit better group fairness. Next, we validate our theoretical findings by conducting extensive experiments on three publicly available datasets-MS-COCO [27], CelebA [31], and FairFace [16] across three different tasks: multi-label classification, facial expression recognition, and gender classification. We carefully evaluate fairness using three widely adopted metrics: Equalized Odds, Equal Opportunity, and Demographic Parity. The results are strongly aligned with our theoretical analysis, consistently demonstrating that adaptive gradient methods achieve better fairness scores compared to stochastic gradient methods.
Our contributions in this paper are summarized as follows: (1) First, we provide a detailed mathematical analysis to study the impact of optimization algorithms on fairness in an analytically tractable setting. Here, we demonstrate that adaptive gradient algorithms like RMSProp have a higher probability of converging to fairer minima. (2) Next, we establish and prove two theorems showing that under appropriate conditions: (a) RMSProp provides fairer updates than SGD, and (b) the worst-case bias introduced by RMSProp in a single optimization step has an upper bound equal to that of SGD. These theoretical insights highlight key properties of optimization algorithms that contribute to their advantages in terms of fairness. (3) Through extensive experiments on multiple publicly available datasets and tasks, we empirically validate our theoretical findings. We systematically evaluate the fairness of models optimized using adaptive gradient methods and compare the outcomes to those obtained from stochastic gradient methods. Our results consistently show that adaptive gradient methods lead to fairer minima. (4) To enable rapid reproducibility and contribute to the area, we make our code publicly available at: https://github.com/Mkolahdoozi/Some-Optimizers-Are-More-Equal.
@@ 2 Related work
Group fairness. Bias in machine learning models leads to discriminatory outcomes, and as these models are increasingly being deployed in real-life applications, such biases can impose risks to society [38]. This risk is particularly important in high-stakes applications, such as hiring [17], healthcare [22], etc. When models yield biased outcomes, they favor certain demographic groups, which reinforces social inequalities. As an example, facial expression recognition models show higher error rates for darker skin tones [55]. Such biases undermine trust in machine learning models and pose ethical and legal challenges. To analyze fairness in machine learning, researchers in this area often distinguish the notion of “group fairness” and “individual fairness” [38]. Group fairness focuses on the notion that machine learning models treat different demographic groups, such as those defined by race, the same. However, individual fairness dictates that similar individuals should be treated similarly. In this paper, we focus on group fairness, which is more widely recognized in practical applications [7].
@@
One of the main sources of bias in machine learning models is data imbalance [9]. When certain demographic subgroups are underrepresented, models prioritize patterns from the majority group, leading to biased decisions. To tackle this, a variety of methods have been proposed that can be categorized into 3 main groups [6]: pre-processing, in-processing, and post-processing. Pre-processing methods often apply some type of transformation to data to make the demographic information less recognizable to the model. As an example, [48] augments the training set with synthetically generated samples, thereby ensuring an equal distribution for demographic subgroups. In-processing methods directly modify the learning algorithm of the model. As an example, [20] proposes a new loss function based on maximum mean discrepancy statistical distance to enhance fairness. Post-processing methods intervene after training is completed. For instance,  [44] introduces a post-processing approach using graph Laplacian regularization to enforce individual fairness constraints while preserving accuracy. Nevertheless, in practice, these fairness-enhancing methods are rarely used in real-world pipelines [39]. One major limitation is their computational overhead or interference with the training phase. As an example, pre-processing methods require extensive data augmentation or transformations. Considering these challenges, an alternative approach to enhance fairness is to use components that are inherently present in every training pipeline. One of the fundamental building blocks of modern deep learning is optimization [52]. Understanding the role of different optimization algorithms in fairness is crucial for practitioners seeking to develop fairer models without additional fairness-specific modifications. The work in [29] provides a key theoretical analysis, showing that standard and unconstrained machine learning implicitly favors one fairness criterion over others. The authors prove that a model’s deviation from calibration is bounded by its excess risk. This is related to our work as it also investigates the fairness properties that emerge from a standard training process. Furthermore, [56] focuses on in-processing fairness methods that use surrogate functions to ensure fairness. Their theory also highlights that balanced datasets are beneficial for achieving tighter fairness and stability guarantees. This is relevant to our findings, as we also identify data imbalance as a critical factor through stochastic differential equation analysis.
@@
Optimization algorithms. Optimization algorithms in deep learning generally fall into two main categories: SGD and adaptive gradient methods (e.g., Adam [18] and RMSProp [34]). SGD updates model parameters using a constant learning rate [15], while adaptive methods dynamically adjust per-parameter learning rates based on past gradients, often leading to faster convergence [35]. The impact of these two types of optimization algorithms on group fairness remains not well understood. Nonetheless, several studies have explored their effects in other areas. For instance, [34] empirically demonstrates that models trained with SGD are more robust to input perturbations than those trained with adaptive methods. Fairness and robustness are often competing objectives in deep learning [36]. This hints that adaptive optimizers may yield better fairness compared to SGD. However, this relationship remains largely unexplored.
@@ 3.1 Preliminaries
Optimization algorithms. Stochastic gradient methods like SGD and adaptive gradient methods such as RMSProp and Adam are widely used for optimizing deep neural networks [58]. Given a loss function $\mathcal{L}(w)$, SGD updates the model parameters $w$ at each iteration $k$ as:
@@
where $\eta$ denotes the learning rate. Adaptive gradient methods improve the speed of convergence by applying coordinate-wise scaling to the gradient [41]. Specifically, RMSProp normalizes the gradient with an exponentially decaying average of its squared values:
@@
Stochastic Differential Equations (SDE). Given the interval $[0,T]$ where $T>0$, we have  [51]:
@@
where $X_{t}\in R^{d}$ is a stochastic process (the solution of the SDE), $a:R^{d}\times[0,T]\rightarrow R^{d}$ is the drift term, $b:R^{d}\times[0,T]\rightarrow R^{d\times s}$ is the diffusion term, and $B(t)\in R^{s}$ is $s$-dimensional Brownian motion. Intuitively, SDEs extend ordinary differential equations (ODEs) by incorporating stochastic noise. Analyzing functions involving stochastic processes requires tools beyond classical calculus. A fundamental theorem in stochastic analysis is Itô’s lemma, extending the chain rule to the functions of stochastic processes [1]. Please see the Appendix A for Itô’s lemma. Unlike ODEs, the solution of an SDE is a stochastic process $X_{t}$. Formally, the solution can be represented as:
$X_{t}=X_{0}+\int_{0}^{t}a(X_{\tau},\tau)d\tau+\int_{0}^{t}b(X_{\tau},\tau)dB_{\tau}.$ (8)
@@
$\int_{0}^{t}b(X_{\tau},\tau)dB_{\tau}=\lim_{n\rightarrow\infty}\sum_{\tau_{k},\tau_{k-1}\in\pi(n)}b(X_{\tau_{k-1}},\tau_{k-1})(B_{\tau_{k}}-B_{\tau_{k-1}}),$ (9)
where $\pi(n)$ denotes a sequence of $n$ partitions in $[0,t]$, and $\lim$ shows convergence in probability. In many cases, explicitly computing $X_{t}$ is intractable. Instead, it is often more practical to analyze the probability distribution of $X_{t}$. The Fokker-Planck equation, also known as the forward Kolmogorov equation, describes the time evolution of the probability density function $p(x,t)$ of $X_{t}$ [21] as:
@@
where $\nabla\cdot$ represents the divergence operator. For completeness, we provide the derivation of the Fokker-Planck equation from Eq. 7 in Appendix B.
SDE approximations for optimization algorithms. SDEs provide a powerful framework for analyzing the dynamics of discrete optimization algorithms such as SGD and RMSProp in a continuous setup. While direct analysis of discrete updates is often challenging [26], SDE approximations enable the study of optimization trajectories. In the context of optimization, the sources of stochasticity include mini-batch gradient noise which are variations due to data heterogeneity (e.g., differences in mini-batches drawn from different subpopulations), among others. To formalize this, we adopt the Noisy Gradient Oracle with Scale Parameter (NGOS) [37], which models stochastic gradients as $g(w)=\nabla\mathcal{L}(w)+\Theta z$, where $\Theta$ is a noise scale parameter, and $z$ is a random variable that follows a distribution with covariance matrix $\Sigma(w)$ and mean of zero. NGOS formally models the gradients used in mini-batch training. It describes the gradient from a small data batch as the “true” gradient (which would come from the full dataset) plus random noise introduced by the sampling process. A well-behaved NGOS is defined as $\nabla\mathcal{L}(w)$ being Lipschitz and $\sqrt{\Sigma(w)}$ being bounded. See Appendix C for a detailed description of a well-behaved NGOS. Note that these assumptions are common and standard in prior works that study optimization using SDEs [25].
Under well-behaved NGOS, SGD can be approximated by the following SDE [25]:
@@
where $W_{t}$ represents an approximation of $w_{k}$ at discrete time steps $k\,\eta$. More recently, SDE approximation for RMSProp has been proposed as a coupled system of equations [37]:
@@ 3.2 SDE analysis of optimizers
where the weights correspond to the proportion of each subgroup in the population. In this setup, the optimal solution for subgroup 0 is $w_{0}^{*}=1$, while the optimal solution for subgroup 1 is $w_{1}^{*}=-1$. However, The population’s optimal solution is obtained by minimizing $\mathcal{L}_{\text{pop}}(w)$, leading to $w_{\text{pop}}^{*}=0$. The following lemma shows that under the demographic parity definition of group fairness [38], $w_{\text{pop}}^{*}=0$ is the fairest minima.
@@
In practice, we do not have access to the full population above, rather we have access to the set $\Omega$ of $N$ individual samples, sampled uniformly from the population. Additionally, set $\Omega$ may not contain an equal number of samples from subgroups 0 and 1, thus introducing bias. We assume that the probability of subgroup 0 in $\Omega$ is $p_{0}$ while the probability of subgroup 1 is $p_{1}=1-p_{0}$. For a given set $\Omega$, the empirical minimization problem used in practice is defined as: $\mathcal{L}_{emp}(w)=\frac{1}{N}\sum_{r\in\Omega}\mathcal{L}_{q_{r}}(w),$ where $q_{r}\in\{0,1\}$ corresponds to samples from subgroups 0 and 1. Note that as $N\rightarrow\infty$, $\mathcal{L}_{emp}\rightarrow\mathcal{L}_{pop}$. In min-batch training with batch size of 1, at each iteration, the optimizer processes a single individual from subgroup 0 with a probability of $p_{0}$ and from subgroup 1 with a probability $p_{1}$. In other words, with both SGD and RMSProp, the model receives $\nabla\mathcal{L}_{0}(w)$ with a probability of $p_{0}$ and $\nabla\mathcal{L}_{1}(w)$ with a probability of $p_{1}$, updating $w$ according to Eqs. 1 and 2, respectively. If $p_{0}>p_{1}$, the optimization trajectory is biased towards the group-specific minimum at $w=1$, and w=1w=-1 otherwise. This demonstrates that disproportionate sampling of subgroups leads to unfair minima, which is particularly relevant in practical settings. The following theorem establishes that when the sampling bias $|p_{0}-p_{1}|$ exceeds a certain threshold, RMSProp has a higher probability of converging to $w^{*}_{pop}$ compared to SGD.
@@
Here, $\epsilon$ is added for numerical stability, but it can be ignored as it is commonly done so in prior work [37]. Next, let’s derive the SDE for SGD:
@@
where $\vartheta=\frac{1}{8\eta p_{0}p_{1}}$. As we use a batch size of 1, we can set $\Theta=1$ [37]. Then, $1>=2\sqrt{p_{0}p_{1}}$ inequality holds. With this in mind, by substituting $w^{*}_{pop}=0$ in Eqs. 20 and 21, we can conclude $\frac{p_{rms}(w^{*}_{pop})}{p_{sgd}(w^{*}_{pop})}>1$ holds if and only if $|p_{0}-p_{1}|>\Delta(p_{1}p_{2},\eta)$, where $\Delta(p_{1}p_{2},\eta)=\sqrt{\frac{\frac{1}{2}\ln{\frac{\vartheta}{\kappa}}}{\vartheta-\kappa}}$. ∎
@@
Now, let’s perform a quick simulation to verify our findings. We use the loss functions defined in Eq. 13 and set $p_{0}=0.1$ and $p_{1}=0.9$ to induce a severe sampling bias. We fix the learning rate to $\eta=0.1$ and apply SGD and RMSProp for 100 epochs. Each experiment is repeated 1000 times, and at every epoch we record the fraction of runs converging to the neighborhood of the fair optimum $w^{*}_{pop}$, defined by $|w-w^{*}_{pop}|<0.2$. The result is shown in Fig. 1 (left). As it is evident from this figure, approximately 10% of the runs under RMSProp lie in the neighborhood of $w^{*}_{pop}$, whereas SGD rarely converges to this fair region. Now, let’s induce a milder bias by setting $p_{0}=0.3$ and $p_{1}=0.7$ and run the same experiment. The results are illustrated in Fig. 1 (right). As shown in this figure, in the milder bias, the two optimizers behave more similarly, yet RMSProp attains a slightly higher fraction of solutions near $w^{*}_{pop}$. These observations align with our theoretical findings in Theorem 1. To show that our findings are true for a wider range of $\eta$ and alternative definitions of the fair neighborhood, we refer the reader to Appendix F for additional experiments.